OpenStudy (anonymous):

HELP! If cos 0 = 5/13 what is the value of sin 0 in quadrant 1?

TheSmartOne (thesmartone):

sin 0 or sin \(\theta\) ?

TheSmartOne (thesmartone):



TheSmartOne (thesmartone):

what is sin in terms of the sides? is it adjacent over hypotenuse? opposite over adjacent? opposite over hypotenuse?

OpenStudy (anonymous):

The 0 with a line in it

TheSmartOne (thesmartone):

what I thought because then sin 0 = 0 :p

OpenStudy (anonymous):

Sorry. I didn't know how to add the line in the middle

TheSmartOne (thesmartone):

`\(\theta\)`

TheSmartOne (thesmartone):

anyways, since cos \(\theta\) = \(\Large \frac{opposite}{hypotenuse}=\frac{5}{13}\) I was able to draw the triangle above.

TheSmartOne (thesmartone):

\(\color{blue}{\text{Originally Posted by}}\) @TheSmartOne what is sin in terms of the sides? is it adjacent over hypotenuse? opposite over adjacent? opposite over hypotenuse? \(\color{blue}{\text{End of Quote}}\) so answer this question and you'll be on your way to your answer :)